# MMGRec: Multimodal Generative Recommendation with Transformer Model

### v2

@@ 2.1. Multimodal Recommendation
Research on multimodal recommendation aims to enhance recommender systems by leveraging multimodal information (Liu et al., 2023a). Existing methodologies typically employ a combination of CF and content-based recommendation techniques (Liu et al., 2023b). Early studies aim to integrate multimodal content for more rational representation learning. For example, the pioneering work VBPR (He and McAuley, 2016) maps visual information to the representation space and concatenates it with CF representations. Subsequent work like ACF (Chen et al., 2017) and UVCAN (Liu et al., 2019) adaptively select multimodal features using different levels of attention mechanisms to learn representations of users and items. In recent years, the research trend turns to exploring the relationship between modal information and users/items. For instance, the representative work MMGCN (Wei et al., 2019) introduces a modality-aware GCN into multimodal recommendation tasks to aggregate and propagate multimodal information on the user-item graph. HHFAN (Cai et al., 2021) utilizes a modality-aware heterogeneous information graph to explore more complex relationships among users, micro-videos, and multimodal information, thereby improving recommendation performance.
@@ 3. Preliminary
With the obtained semantic ID, a Transformer-based sequence-to-sequence model is trained on the sequences of semantic IDs corresponding to the items from a user’s chronological interactions. The objective is to predict the semantic ID of the next item after the sequence. This model autoregressively decodes the tokens of semantic ID identifying the next item, thereby qualifying as a generative recommendation.
@@ 5.2. Baselines
  • NGCF (Wang et al., 2019) encodes CF signals into representation learning by exploiting high-order connectivity from user-item interactions.
@@
  • VBPR (He and McAuley, 2016) incorporates multimodal information with MF framework to predict the interactions between users and items.
  • MMGCN (Wei et al., 2019) allocates an independent GCN for each modality, which learns modality-specific user preferences and item representations via the propagation of modality information.
@@
  • LATTICE (Zhang et al., 2021) discovers item relationships via multimodal features to learn a graph structure. Graph convolutions aggregate high-order affinities along the structure for recommendation.
@@
  • LightGT (Wei et al., 2023) is a state-of-the-art Transformer-based model. It designs modal-specific embedding and layer-wise position encoding for effective feature distillation, learning superior user preference on item content for interaction prediction.
@@ 6.1. Performance Comparison (RQ1)
Table 2 presents the results of MMGRec and baselines over three experimental datasets. Besides, it reports the improvements and statistical significance test, which are calculated between our method and the strongest baseline (highlighted with underline). From the results, the main observations are as follows:
  • MMGRec consistently achieves the best performance in all cases. In particular, improvements over the strongest competitor w.r.t. NDCG@10 are 7.17%, 6.79%, and 6.58% in MovieLens, TikTok, and Kwai, respectively. Additionally, we conduct one-sample t-tests, which reveal that the improvements of MMGRec are statistically significant (p-value << 0.05). The primary difference between MMGRec and baselines lies in whether the recommendation paradigm is generative or traditional. Hence, we attribute the significant improvements to introducing a Transformer-based generative paradigm in multimodal recommendation.
@@ 6.5. Analysis of Rec-ID Design (RQ5)
Impact of Rec-ID Collision. As shown in Figure 6(a), the bar plots show the collision rate on the Kwai dataset under different data scales, measured as the proportion of items assigned non-unique Rec-IDs. The x-axis value 1/16 indicates a subset sampled from one-sixteenth of the items and their interacting users. As the dataset scale increases, the collision rate rises, and model performance degrades slightly, indicating that Rec-ID collisions have a non-negligible impact in large-scale recommendation scenarios. The line plots compare our popularity token with a random token (as in TIGER). With higher collision rates, the popularity token achieves larger performance gains. This indicates that popularity-aware semantic ordering better distinguishes colliding Rec-IDs and reduces the performance degradation caused by collisions.
Effect of Rec-ID Length. Figure 6(b) shows the performance trends in Recall@10 and NDCG@10 as the Rec-ID length $M$ increases from 2 to 5. To cover the full item set, a shorter Rec-ID results in a larger per-step decoding search space. As $M$ increases from 2 to 4, model performance steadily improves, since overly short Rec-IDs increase decoding difficulty due to a large set of candidate tokens at each step. When $M=4$, the model achieves a good balance between the number of decoding steps and the per-step search space, leading to the best performance. However, when $M$ further increases to 5, more autoregressive steps introduce higher generation difficulty and inference latency, resulting in performance degradation. Therefore, we set $M=4$ by default.

### v3

@@ 2.1. Multimodal Recommendation
Research on multimodal recommendation aims to enhance recommender systems by leveraging multimodal information (Liu et al., 2023a). Existing methodologies typically employ a combination of CF and content-based recommendation techniques (Liu et al., 2023b). Early studies aim to integrate multimodal content for more rational representation learning. For example, the pioneering work VBPR (He and McAuley, 2016) maps visual information to the representation space and concatenates it with CF representations. Subsequent work like ACF (Chen et al., 2017) and UVCAN (Liu et al., 2019) adaptively select multimodal features using different levels of attention mechanisms to learn representations of users and items. In recent years, the research trend turns to exploring the relationship between modal information and users/items. For instance, the representative work MMGCN (Wei et al., 2019) introduces a modality-aware GCN into multimodal recommendation tasks to aggregate and propagate multimodal information on the user-item graph. HHFAN (Cai et al., 2021) constructs a modality-aware heterogeneous graph to model intricate interactions among users, micro-videos, and multimodal signals, thereby improving recommendation performance.
@@ 3. Preliminary
With the obtained semantic ID, a Transformer-based sequence-to-sequence model is trained on sequences of semantic IDs derived from a user’s chronological item interactions. The objective is to predict the semantic ID of the next item after the sequence. This model autoregressively decodes the tokens of semantic ID identifying the next item, thereby qualifying as a generative recommendation.
@@ 5.2. Baselines
  • NGCF (Wang et al., 2019) encodes CF signals into representation learning by leveraging high-order connectivity from user-item interactions.
@@
  • VBPR (He and McAuley, 2016) integrates multimodal information into the MF framework to predict user-item interactions.
  • MMGCN (Wei et al., 2019) assigns a dedicated GCN to each modality to learn modality-specific user preferences and item representations through modality-wise information propagation.
@@
  • LATTICE (Zhang et al., 2021) discovers item relationships via multimodal features to construct a graph structure, on which graph convolutions aggregate high-order affinities for recommendation.
@@
  • LightGT (Wei et al., 2023) is a state-of-the-art Transformer-based model. It designs modal-specific embedding and layer-wise position encoding for effective feature distillation, learning high-quality user preferences on item content for interaction prediction.
@@ 6.1. Performance Comparison (RQ1)
Table 2 summarizes the performance of MMGRec and the baseline methods on three experimental datasets. It also reports the relative improvements and statistical significance tests, computed between our model and the strongest baseline (underlined). From these results, we make the following observations:
  • MMGRec consistently delivers the best performance across all settings. In particular, it outperforms the strongest competitor in terms of NDCG@10 by 7.17%, 6.79%, and 6.58% on MovieLens, TikTok, and Kwai, respectively. We further perform one-sample t-tests, which indicate that the gains achieved by MMGRec are statistically significant (p-value << 0.05). The key distinction between MMGRec and the baseline methods lies in the adoption of a generative rather than a conventional recommendation paradigm. Therefore, we attribute the substantial performance improvements to the introduction of a Transformer-based generative framework for multimodal recommendation.
@@ 6.5. Analysis of Rec-ID Design (RQ5)
Impact of Rec-ID Collision. As shown in Figure 6(a), the bar plots show the collision rate on the Kwai dataset under different data scales, measured as the percentage of items assigned non-unique Rec-IDs. The x-axis value 1/16 indicates a subset sampled from one-sixteenth of the items and their interacting users. As the dataset scale increases, the collision rate rises, and model performance degrades slightly, indicating that Rec-ID collisions have a non-negligible impact in large-scale recommendation scenarios. The line plots compare our popularity token with a random token (as in TIGER). With increasing collision rates, the popularity token outperforms the random token by a larger margin. This indicates that popularity-aware semantic ordering better distinguishes colliding Rec-IDs and reduces the performance degradation caused by collisions.
Effect of Rec-ID Length. Figure 6(b) shows the performance trends in Recall@10 and NDCG@10 as the Rec-ID length $M$ increases from 2 to 5. To cover the full item set, a shorter Rec-ID results in a larger per-step decoding search space. As $M$ increases from 2 to 4, model performance steadily improves, since overly short Rec-IDs increase decoding difficulty due to a large set of candidate tokens at each step. When $M=4$, the model achieves a good balance between the number of decoding steps and the per-step search space, leading to the best performance. However, when $M$ further increases to 5, more autoregressive steps introduce higher generation difficulty and inference latency, resulting in performance degradation (Wang et al., 2024). Therefore, we set $M=4$ by default.